# Separation and collapse of equilibria inequalities on AND–OR trees without shape constraints

### v2

@@ Abstract
Herein, we investigate the zero-error randomized complexity, which is the least cost against the worst input, of AND–OR tree computation by imposing various restrictions on the algorithm to find the Boolean value of the root of that tree and no restrictions on the tree shape. When a tree satisfies a certain condition regarding its symmetry, directional algorithms proposed by Saks and Wigderson (1986), special randomized algorithms, are known to achieve the randomized complexity. Furthermore, there is a known example of a tree that is so unbalanced that no directional algorithm achieves the randomized complexity (Vereshchagin 1998). In this study, we aim to identify where deviations arise between the general randomized Boolean decision tree and its special case, directional algorithms. In this paper, we show that for any AND–OR tree, randomized depth-first algorithms, which form a broader class compared with directional algorithms, have the same equilibrium as that of the directional algorithms. Thus, we get the collapse result on equilibria inequalities that holds for an arbitrary AND–OR tree. This implies that there exists a case where even depth-first algorithms cannot be the fastest, leading to the separation result on equilibria inequality. Additionally, a new algorithm is introduced as a key concept for proof of the separation result.
@@ 1.1 Background
AND–OR tree is a computation model of a read-once Boolean function, that is, a function in which each of its variable appears exactly once in its formula. We are interested in the decision tree complexity (or deterministic complexity) of a tree here, i.e., the computational cost is measured by the number of leaves the algorithm (i.e., the decision tree) has queried. $\alpha$ runs the algorithms computing the tree and $\omega$ runs the assignments to the tree, then the decision tree complexity is defined as follows:
@@
It is well known that the randomized complexity may be smaller than the deterministic complexity [14]. In other words, randomization may improve efficiency. There is a considerable amount of prior research on separation between deterministic complexity and randomized complexity. In particular, an open problem in [14] has been solved in mid 2010s, and remarkable developments continue to take place [2, 10, 9].
Note the differences in symbols in the literature. Our definition of $R$ is the same as [14]. We later define $R_{0}$ and $R_{1}$ also in the style of [14]; let us call this notation “classical style”. On the other hand, in some recent literature [2, 10, 9], the symbols $R$, $R_{0}$, and $R_{1}$ are used to mean something different from what they do in [14]; let us call it “modern style”. Table 1 provides summary of differences. In the row of “zero error, the root is $i$”, the truth assignment is restricted to those with root value $i$, and the algorithm is zero-error regardless of the root value. This paper adopt classical style.
@@
Table 1: Symbols for variants of randomized complexity
It is reasonable to restrict ourselves to algorithms with alpha–beta pruning procedures: For each OR node $v$, we assume that once an algorithm finds a child node of $v$ has value 1 (true) then the algorithm does not search the rest of the descendants of $v$ further and finds that $v$ has value 1, and we make similar assumptions for AND nodes. We shall review this topic in Definition 2.2.
It is often the case that depth-first algorithms are considered to examine essential instances without any unnecessary complication. A randomized algorithm of each type (general or depth-first) is defined as a probability distribution on each class of algorithms. Moreover, we consider a special type of randomized algorithms called directional algorithms proposed by Saks and Wigderson [14]. We refer to directional algorithms in this sense as RDA for technical reasons, which stands for randomized directional algorithm. All these types of algorithms will be precisely defined in Section 2. Inclusion relationships between algorithm classes are as follows:
[…]
For a fixed tree $T$, by restricting the type of randomized algorithms in (1.2) to randomized depth-first algorithms (RDAs, repectively), we define $R_{\mathrm{DF}}(T)$ ($d(T)$, respectively). Precise definitions of $R_{\mathrm{DF}}(T)$ and $d(T)$ are given in Definition 2.5. We often consider dual complexity $P_{\mathrm{DF}}(T)$ by restricting the algorithm type in (1.3) to depth-first ones. Note that a set of all randomized depth-first algorithms is convex in the following sense. For any randomized depth-first algorithms $A$ and $B$ and a nonnegative real number $p\leq 1$, the algorithm $pA+(1-p)B$ (the algorithm that works as $A$ with probability $p$ and as $B$ with probability $1-p$) is also a randomized depth-first algorithm. Using the properties of convex sets, we obtain an important property that the randomized complexity and its dual are equal.
Yao’s principle [20]: For any AND–OR tree $T$, we have $R(T)=P(T)$, and $R_{\mathrm{DF}}(T)=P_{\mathrm{DF}}(T)$ (We will review the proof for this in Proposition 2.7).
While RDA has the advantage of being easy to use via induction on tree height, it also has the drawback of not being convex. Thus, Yao’s principle is not applicable for RDAs. We can define a minimum convex set of randomized algorithms containing all RDAs, which lies between the randomized depth-first algorithms and RDAs in the inclusion relationships (1.4). We let “a randomized directional algorithm in the broad sense” to denote an element of this convex set. A detailed definition of a randomized directional algorithm in the broad sense is given in Definition 2.2. $R_{\mathrm{dir}}(T)$ and $P_{\mathrm{dir}}(T)$ are defined in the same way as above for randomized directional algorithms in the broad sense, and the Yao’s principle holds in this case: $R_{\mathrm{dir}}(T)=P_{\mathrm{dir}}(T)$.
Previous studies have reported various results regarding algorithms of the above-mentioned types. A balanced tree is a tree such that the internal nodes of the same depth (distance from the root) have the same number of child nodes, and all leaves have the same depth. In [8, 16, 11], it was shown that for every complete binary tree (or more generally, a balanced tree), there exists a unique randomized assignment that attains $P(T)$ (the same applies to $P_{\mathrm{DF}}(T)$), but infinitely many assignments attains $P_{\mathrm{dir}}(T)$. [17], [12], and [13] investigated the case in which randomized assignments are restricted to ones with independent probabilities for each leaves, and showed that directional algorithms are sufficient to achieve the minimum cost for such assignments.
[14] and [7] revealed some important facts related to our result; when we consider a tree with certain mild condition applied to its symmetry (which we call the weak-balance condition), $R=d$ holds.
@@
[7] (implicitly in [14]) Let $T$ be an AND–OR tree. For each node $v$ of $T$, let $T_{v}$ be the subtree of $T$ where the root node is $v$, and let $R_{i}(T_{v})$ be the randomized complexity of $T_{v}$ with respect to the assignments setting the value of $v$ to $i$. $T$ is said to be weakly balanced if the following holds: For each internal node $v$ of $T$ and its children $v_{1},\ldots,v_{n}$, if $v$ is an AND (OR, respectively) node, then $R_{0}(T_{v_{j}})\leq R_{0}(T_{v_{k}})+R_{1}(T_{v_{j}})$ ($R_{1}(T_{v_{j}})\leq R_{1}(T_{v_{k}})+R_{0}(T_{v_{j}})$, respectively) for each $j,k\in\{1,\ldots,n\}$.
It is known that every complete binary tree and balanced tree is weakly balanced. The following fact is significant. A more concise proof is provided in [4].
Theorem 1.2.
[7, Theorem 2.1] (see also [14, Lemma 5.1]) Suppose an AND–OR tree $T$ is weakly balanced. Then $R(T)=d(T)$.
Example 1.3.
[19] (see also [1]) There exists an AND–OR tree $V$ where
$R(V)\leq 51,\ d(V)=\frac{33525}{640}=52.38\ldots$
1.2 Our goal: Collapse and separation
Our main motivations for focusing on the directional algorithms is precisely because Example 1.3 holds. Considering Example 1.3, it is natural to ask where does the gap lies between $R$ and $d$. We show a collapse on equilibria
$R_{\mathrm{DF}}(T)=d(T)$ (1.5)
for all AND–OR tree $T$: The randomized complexity with respect to the depth-first algorithms equals that with respect to the directional algorithms (RDAs). Here, no shape constraints are imposed on the tree $T$. Any node of $T$ can have any number (${\geq}2$) of child nodes, and there can be a mixture of AND and OR nodes among the child nodes. (1.5) is a strict equality without Landau symbol $O$. Thus, we obtain a separation result on the equilibria
$R(T)<R_{\mathrm{DF}}(T)$ (1.6)
1.3 Our method: The “chimera” algorithm
The ideal scenario for proof is as follows.
(i) Consider an algorithm $A$ that achieves $R_{\mathrm{DF}}(T)$.
(ii) By adopting a certain induction hypothesis, it is shown that there exists a good algorithm $B_{i}$ for each subtree $T_{i}$ just under the root of $T$.
(iii) Replace the behavior of $A$ in subtree $T_{i}$ with algorithm $B_{i}$ for each $i$. Then, let $A^{\prime}$ be the resulting algorithm.
(iv) We want to show that $A^{\prime}$ is an RDA, and that the cost of $A^{\prime}$ is lower than or equal to the cost of $A$.
(v) We prove that $R_{\mathrm{DF}}(T)=d(T)$ using Yao’s principle if necessary.
However, when it is time to implement this scenario in a proof, the following difficulties arise: First, when $A$ in (i) is not an RDA, the meaning of the operations described in (iii) is unclear. Second, as mentioned earlier, we cannot apply the Yao’s principle to RDAs.
To overcome these difficulties, we divide the problem of showing $R_{\mathrm{DF}}(T)=d(T)$ into two problems, namely, Problem 1: $R_{\mathrm{dir}}(T)=d(T)$ and Problem 2: $R_{\mathrm{DF}}(T)=R_{\mathrm{dir}}(T)$. In Section 3, we will solve Problem 1.
Problem 2 is more difficult to solve. In Section 4, to solve Problem 2, we investigate the case when $A$ in (i) is a deterministic depth-first algorithm $\alpha$. Instead of $A^{\prime}$ in (iii), we elaborately define a randomized directional algorithm in the broad sense $B_{\alpha,\delta}$, which is a key concept in this study. Figure 1 in subsection 4.1 illustrates interdependence among concepts necessary to define $B_{\alpha,\delta}$. The nickname “chimera” for algorithm $B_{\alpha,\delta}$ arises from the fact that $B_{\alpha,\delta}$ is composed of parts from several other algorithms and probability distributions as well as $\alpha$.
In Section 2, we give the basic definitions and list some fundamental facts regarding equilibrium values. In Section 5, we summarize our results and describe some immediate consequences along with those reported in previous studies. In Section 6, we provide rigorous proofs of some facts used in other sections.
2 Definitions and Preliminaries
In this section, we will define some basic concepts and see some fundamental facts on AND–OR trees. See also [3, Chapter 12] and [15] for helpful examples.
Let $T$ be an AND–OR tree whose root has $n$ child nodes $v_{1},\dots,v_{n}$. Let $T_{1},\dots,T_{n}$ be the subtrees whose roots are $v_{1},\dots,v_{n}$ respectively.
Convention 2.1.
Throughout this paper, no restrictions are placed on the shape of an AND–OR tree unless otherwise specified. An internal node may have arbitrary number ($\geq 2$) of child nodes. We also relax the constraints that the tree is alternating; to be more precise, each internal node of an AND–OR tree $T$ may be labeled with AND or OR regardless of its position, in other words, regardless of whether it is a root or not, and regardless of whether its parent node is labeled with AND or OR. $T$ may not be weakly balanced (Definition 1.1).
Definition 2.2.
Suppose that $T$ is as above.
@@ 2 Definitions and Preliminaries
    (Assignment) The value of the root of $T$ is a Boolean function of the values of the leaves of $T$. An assignment to $T$ is an $\{0,1\}$ string of such input values to the leaves. More precisely, if $\ell_{1},\ldots,\ell_{n}$ are the leaves of $T$, then an assignment is a mapping $\omega:\{\ell_{1},\ldots,\ell_{n}\}\rightarrow\{0,1\}$.
@@
    ($\mathcal{A}_{\mathrm{dir}}(T)$, the directional algorithms) A deterministic depth-first algorithm is said to be directional if the algorithm has a fixed order of leaf queries, and does not skip the queries unless in the case where it omits by alpha–beta pruning procedure explained above. Note that in this paper, a directional algorithm is always depth-first. We let $\mathcal{A}_{\mathrm{dir}}(T)$ denote the set of all deterministic directional algorithms on $T$. Thus we have $\mathcal{A}_{\mathrm{dir}}(T)\subseteq\mathcal{A}_{\mathrm{DF}}(T)\subseteq\mathcal{A}(T)$. It is easily seen that the two subset symbols are proper.
@@
    A randomized algorithm on $T$ is an element of $\mathcal{D}(\mathcal{A}(T))$, that is, a probability distribution on $\mathcal{A}(T)$. For a randomized algorithm $A$ on $T$ and each $\alpha\in\mathcal{A}(T)$, $A(\alpha)$ denotes the probability of $\alpha$.
@@
    A randomized depth-first algorithms is an element of $\mathcal{D}(\mathcal{A}_{\mathrm{DF}}(T))$, that is, a probability distribution on $\mathcal{A}_{\mathrm{DF}}(T)$. A randomized directional algorithm in the broad sense is an element of $\mathcal{D}(\mathcal{A}_{\mathrm{dir}}(T))$, that is, a probability distribution on $\mathcal{A}_{\mathrm{dir}}(T)$.
@@
    (RDA) A randomized directional algorithm in the sense of Saks and Wigderson (RDA, for short) is recursively defined. A randomized algorithm $A$ is an RDA if there exist a sequence of RDA $\langle A_{1},\dots,A_{n}\rangle$ on the subtrees $T_{1},\dots,T_{n}$ and a probability distribution $\pi$ on $n$th symmetric group $\mathfrak{S}_{n}$ with the following property. For each permutation $\sigma$ on $\{1,\dots,n\}$, with probability $\pi(\sigma)$, $A$ performs $A_{\sigma(1)},\dots,A_{\sigma(n)}$ in this order.
@@
For the above-mentioned tree $T$, the set of all alpha–beta pruning algorithms in the sense of Knuth and Moore (1975) [6] equals $\mathcal{A}_{\mathrm{DF}}(T)$.
Saks and Wigderson [14] defined RDA in order to get a recurrence formula of randomized complexity. In general, a probability distribution on the deterministic directional algorithms is not necessarily an RDA .

### v3

@@ Abstract
Herein, we investigate the zero-error randomized complexity, which is the least cost against the worst input, of AND–OR tree computation by imposing various restrictions on the algorithm to find the Boolean value of the root of that tree and no restrictions on the tree shape. When a tree satisfies a certain condition regarding its symmetry, directional algorithms proposed by Saks and Wigderson (1986), special randomized algorithms, are known to achieve the randomized complexity. Furthermore, there is a known example of a tree that is so unbalanced that no directional algorithm achieves the randomized complexity (Vereshchagin 1998). In this study, we aim to identify where deviations arise between the general randomized Boolean decision tree and its special case, directional algorithms. We show that for any AND–OR tree, randomized depth-first algorithms, which form a broader class compared with directional algorithms, have the same equilibrium as that of the directional algorithms. Thus, we get the collapse result on equilibria inequalities that holds for an arbitrary AND–OR tree. This implies that there exists a case where even depth-first algorithms cannot be the fastest, leading to the separation result on equilibria inequality. Additionally, a new algorithm is introduced as a key concept for proof of the separation result.
@@ 1.1 Background
AND–OR tree is a computation model of a read-once Boolean function, that is, a function in which each of its variables appears exactly once in its formula. In other words, a tree $T$ is an AND-OR tree if each internal node is labeled with AND or OR, and each leaf is a different Boolean variable. We are interested in the decision tree complexity (or deterministic complexity) of a tree here, i.e., the computational cost is measured by the number of leaves the algorithm (i.e., the decision tree) has queried. $\alpha$ runs the algorithms computing the tree and $\omega$ runs the assignments to the tree, then the decision tree complexity is defined as follows:
@@
It is well known that the randomized complexity may be smaller than the deterministic complexity [SW86]. In other words, randomization may improve efficiency. There is a considerable amount of prior research on separation between deterministic complexity and randomized complexity. In particular, an open problem in [SW86] has been solved in mid 2010s, and remarkable developments continue to take place [ABB+17, MS15, MRS18].
Note the differences in symbols in the literature. Our definition of $R$ is the same as [SW86]. We later define $R_{0}$ and $R_{1}$ also in the style of [SW86]; let us call this notation “classical style”. On the other hand, in some recent literature [ABB+17, MS15, MRS18], the symbols $R$, $R_{0}$, and $R_{1}$ are used to mean something different from what they do in [SW86]; let us call it “modern style”. Table 1 provides summary of differences. In the row of “zero error, the root is $i$”, the truth assignment is restricted to those with root value $i$, and the algorithm is zero-error regardless of the root value. This paper adopt classical style.
Table 1: Symbols for variants of randomized complexity
@@
It is reasonable to restrict ourselves to algorithms with alpha–beta pruning procedures: For each OR node $v$, we assume that once an algorithm finds a child node of $v$ has value 1 (true) then the algorithm does not search the rest of the descendants of $v$ further and finds that $v$ has value 1, and we make similar assumptions for AND nodes.
It is often the case that depth-first algorithms are considered to examine essential instances without any unnecessary complication. A randomized algorithm of each type (general or depth-first) is defined as a probability distribution on each class of algorithms. Moreover, we consider a special type of randomized algorithms called directional algorithms proposed by Saks and Wigderson [SW86]. We refer to directional algorithms in this sense as RDA for technical reasons, which stands for randomized directional algorithm. The following is a somewhat lengthy definition, but it is important and should be stated in full here.
Convention 1.1.
Throughout this paper, no restrictions are placed on the shape of an AND–OR tree unless otherwise specified. An internal node may have arbitrary number ($\geq 2$) of child nodes. We also relax the constraints that the tree is alternating; to be more precise, each internal node of an AND–OR tree $T$ may be labeled with AND or OR regardless of its position, in other words, regardless of whether it is a root or not, and regardless of whether its parent node is labeled with AND or OR.
@@
Suppose that $T$ is an AND-OR tree in the sense of Convention 1.1.
@@
    (Assignment) The value of the root of $T$ is a Boolean function of the values of the leaves of $T$. An assignment to $T$ is a $\{0,1\}$ string of such input values to the leaves. More precisely, if $\ell_{1},\ldots,\ell_{n}$ are the leaves of $T$, then an assignment is a mapping $\omega:\{\ell_{1},\ldots,\ell_{n}\}\rightarrow\{0,1\}$.
@@
    ($\mathcal{A}_{\mathrm{dir}}(T)$, the directional algorithms) A deterministic depth-first algorithm is said to be directional if the algorithm has a fixed order of leaf queries, and does not skip the queries unless in the case where it omits by alpha–beta pruning procedure explained above. Note that in this paper, a directional algorithm is always depth-first. We let $\mathcal{A}_{\mathrm{dir}}(T)$ denote the set of all deterministic directional algorithms on $T$. Thus we have $\mathcal{A}_{\mathrm{dir}}(T)\subseteq\mathcal{A}_{\mathrm{DF}}(T)\subseteq\mathcal{A}(T)$. The two subset symbols are proper. The former will be seen later in Example 2.1.
@@
    (Randomized algorithm) A randomized algorithm on $T$ is an element of $\mathcal{D}(\mathcal{A}(T))$, that is, a probability distribution on $\mathcal{A}(T)$. For a randomized algorithm $A$ on $T$ and each $\alpha\in\mathcal{A}(T)$, $A(\alpha)$ denotes the probability of $\alpha$.
@@
    (Randomized depth-first algorithm,) A randomized depth-first algorithms is an element of $\mathcal{D}(\mathcal{A}_{\mathrm{DF}}(T))$, that is, a probability distribution on $\mathcal{A}_{\mathrm{DF}}(T)$.
@@
    (RDA) A randomized directional algorithm in the sense of Saks and Wigderson (RDA, for short) is recursively defined. A randomized algorithm $A$ is an RDA if there exist a sequence of RDA $\langle A_{1},\dots,A_{n}\rangle$ on the subtrees $T_{1},\dots,T_{n}$ just under the root and a probability distribution $\pi$ on $n$th symmetric group $\mathfrak{S}_{n}$ with the following property. For each permutation $\sigma$ on $\{1,\dots,n\}$, with probability $\pi(\sigma)$, $A$ performs $A_{\sigma(1)},\dots,A_{\sigma(n)}$ in this order.
@@
For the above-mentioned tree $T$, the set of all alpha–beta pruning algorithms in the sense of Knuth and Moore (1975) [KM75] equals $\mathcal{A}_{\mathrm{DF}}(T)$.
Inclusion relationships between algorithm classes are as follows:
[…]
For a fixed tree $T$, by restricting the type of randomized algorithms in (1.2) to randomized depth-first algorithms (RDAs, repectively), we define $R_{\mathrm{DF}}(T)$ ($d(T)$, respectively). Precise definitions of $R_{\mathrm{DF}}(T)$ and $d(T)$ are given in Definition 2.3. We often consider dual complexity $P_{\mathrm{DF}}(T)$ by restricting the algorithm type in (1.3) to depth-first ones. Note that the set of all randomized depth-first algorithms is convex in the following sense. For any randomized depth-first algorithms $A$ and $B$ and a nonnegative real number $p\leq 1$, the algorithm $pA+(1-p)B$ (the algorithm that works as $A$ with probability $p$ and as $B$ with probability $1-p$) is also a randomized depth-first algorithm. Using the properties of convex sets, we obtain an important property that the randomized complexity and its dual are equal.
Yao’s principle [Y77]: For any AND–OR tree $T$, we have $R(T)=P(T)$, and $R_{\mathrm{DF}}(T)=P_{\mathrm{DF}}(T)$ (We will review the proof for this in Proposition 2.5).
While RDA has the advantage of being easy to use via induction on tree height, it also has the drawback of not being convex. Thus, Yao’s principle is not applicable for RDAs. We can define a minimum convex set of randomized algorithms containing all RDAs, which lies between the randomized depth-first algorithms and RDAs in the inclusion relationships (1.5). We let “a randomized directional algorithm in the broad sense” to denote an element of this convex set.
Definition 1.4.
(Randomized directional algorithm in the broad sense) A randomized directional algorithm in the broad sense is an element of $\mathcal{D}(\mathcal{A}_{\mathrm{dir}}(T))$, that is, a probability distribution on $\mathcal{A}_{\mathrm{dir}}(T)$.
$R_{\mathrm{dir}}(T)$ and $P_{\mathrm{dir}}(T)$ are defined in the same way as above for randomized directional algorithms in the broad sense, and the Yao’s principle holds in this case: $R_{\mathrm{dir}}(T)=P_{\mathrm{dir}}(T)$.
Previous studies have reported various results regarding algorithms of the above-mentioned types. A balanced tree is a tree such that the internal nodes of the same depth (distance from the root) have the same number of child nodes, and all leaves have the same depth. In [LT07, SN12, POL+16], it was shown that for every complete binary tree (or more generally, a balanced tree), there exists a unique randomized assignment that attains $P(T)$ (the same applies to $P_{\mathrm{DF}}(T)$), but infinitely many assignments attains $P_{\mathrm{dir}}(T)$. [SN15], [PPN+17], and [PPT22] investigated the case in which randomized assignments are restricted to ones with independent probabilities for each leaves, and showed that directional algorithms are sufficient to achieve the minimum cost for such assignments.
[SW86] and [KSS22] revealed some important facts related to our result; when we consider a tree with certain mild condition applied to its symmetry (which we call the weak-balance condition), $R=d$ holds.
Definition 1.5.
[KSS22] (implicitly in [SW86]) Let $T$ be an AND–OR tree. For each node $v$ of $T$, let $T_{v}$ be the subtree of $T$ where the root node is $v$, and let $R_{i}(T_{v})$ be the randomized complexity of $T_{v}$ with respect to the assignments setting the value of $v$ to $i$. $T$ is said to be weakly balanced if the following holds: For each internal node $v$ of $T$ and its children $v_{1},\ldots,v_{n}$, if $v$ is an AND (OR, respectively) node, then $R_{0}(T_{v_{j}})\leq R_{0}(T_{v_{k}})+R_{1}(T_{v_{j}})$ ($R_{1}(T_{v_{j}})\leq R_{1}(T_{v_{k}})+R_{0}(T_{v_{j}})$, respectively) for each $j,k\in\{1,\ldots,n\}$.
It is known that every complete binary tree and balanced tree is weakly balanced. The following fact is significant. A more concise proof is provided in [I24].
Theorem 1.6.
[KSS22, Theorem 2.1] (see also [SW86, Lemma 5.1]) Suppose an AND–OR tree $T$ is weakly balanced. Then $R(T)=d(T)$.
@@
[V98] (see also [A11]) There exists an AND–OR tree $V$ where
$R(V)\leq 51,\ d(V)=\frac{33525}{640}=52.38\ldots$
1.2 Our goal: Collapse and separation
Our main motivations for focusing on the directional algorithms is precisely because Example 1.7 holds. Considering Example 1.7, it is natural to ask where does the gap lies between $R$ and $d$.
Incidentally, in Saks-Wigderson [SW86] and in Vereshchagin [V98], there appears to be an implicit consensus that instead of considering $R_{\mathrm{DF}}$, it is sufficient to consider $d$. This would be based on an intuition that it is not surprising for $R_{\mathrm{DF}}$ and $d$ to be equal.
Observation 1.8.
Suppose that a randomized depth-first algorithm probes subtrees $T_{1}$ and $T_{2}$ under a node $v$ of a tree. The intuitive observation would be that, since probing $T_{1}$ and probing $T_{2}$ are “independent”, when performing depth-first-probing $T_{1}$, there is no need to consider $T_{2}$ in choosing an “optimal” algorithm to probe $T_{1}$, nor is it necessary to consider the history prior to starting the probe of $T_{1}$.
It must be noted that Observation 1.8 is not true in the case where we consider arbitrary depth-first algorithms. The intuition is false because the “independence” fails in the following two ways.
Let $A$ be a randomized depth-first algorithm. Firstly, when we consider arbitrary randomized depth-first algorithms, probabilistic independence between the sub-algorithms of $A$ does not hold by its definition.
Secondly, the sub-algorithms may not be independent in some cases, even if $A$ is deterministic. Suppose that $A$ is deterministic, and it evaluates $T_{1}$ first and then $T_{2}$. As we will see in Example 2.1, It is possible that the query order in $T_{2}$ changes depending on the query result of $T_{1}$. Thus, the movement of a depth-first algorithm in subtrees is not always independent in the sense of query order, either.
Furthermore, there is a more important point that must be noted: The hierarchy of equilibrium values collapses when we make mild assumptions about the symmetry of the tree.
It is immediate that $R(T)\leq R_{\mathrm{DF}}(T)\leq R_{\mathrm{dir}}(T)\leq d(T)$. Under a certain hypothesis on $T$ (weakly balanced in the sense of Definition 1.5), it is shown that $R(T)=d(T)$ (Theorem 1.6).
Various types of trees, including binary trees, are weakly-balanced. So we must ask: In Observation 1.8, have we not unconsciously made assumptions about the shape of the tree?
So we paraphrase the previous question, where does the gap lies between $R$ and $d$, in a more particular form. What assumptions about the symmetry of the tree are sufficient conditions for deriving $R_{\mathrm{DF}}=d$? Our conclusion is that no assumptions need be made.
We show a collapse on equilibria
$R_{\mathrm{DF}}(T)=d(T)$ (1.6)
for all AND–OR tree $T$, regardless of the assumption of weakly balanced. The randomized complexity with respect to the depth-first algorithms equals that with respect to the directional algorithms (RDAs). Here, no shape constraints are imposed on the tree $T$. Any node of $T$ can have any number (${\geq}2$) of child nodes, and there can be a mixture of AND and OR nodes among the child nodes. (1.6) is a strict equality without Landau symbol $O$. Thus, we obtain a separation result on the equilibria
$R(T)<R_{\mathrm{DF}}(T)$ (1.7)
1.3 Technical obstacles
The ideal scenario for proof is as follows.
  1. (i)
    Consider an algorithm $A$ that achieves $R_{\mathrm{DF}}(T)$.
  2. (ii)
    By adopting a certain induction hypothesis, it is shown that there exists a good algorithm $B_{i}$ for each subtree $T_{i}$ just under the root of $T$.
  3. (iii)
    Replace the behavior of $A$ in subtree $T_{i}$ with algorithm $B_{i}$ for each $i$. Then, let $A^{\prime}$ be the resulting algorithm.
  4. (iv)
    We want to show that $A^{\prime}$ is an RDA, and that the cost of $A^{\prime}$ is lower than or equal to the cost of $A$.
  5. (v)
    We prove that $R_{\mathrm{DF}}(T)=d(T)$ using Yao’s principle if necessary.
However, when it is time to implement this scenario in a proof, the following difficulties arise.
(Obstacle 1)
As mentioned earlier, we cannot apply the Yao’s principle to RDAs.
(Obstacle 2)
When $A$ in (i) is not an RDA, the meaning of the operations described in (iii) is unclear.
1.4 Ways to overcome difficulties
Big-picture strategy:   To overcome Obstacle 1, we divide the problem of showing $R_{\mathrm{DF}}(T)=d(T)$ into two problems.
  • Problem 1: $R_{\mathrm{dir}}(T)=d(T)$,
  • Problem 2: $R_{\mathrm{DF}}(T)=R_{\mathrm{dir}}(T)$.
The set of all randomized directional algorithms in the broad sense equals the convex hull of the set of all RDAs. We can apply Yao’s principle to randomized directional algorithms in the broad sense.
Problem 1:   In Section 3, we solve Problem 1. To overcome Obstacle 2 in Problem 1, we provide precise definition on replacement of sub-algorithm. Suppose that $T$ is an AND-OR tree and that $T_{1},\dots,T_{n}$ are subtrees just under the root. Suppose that $\alpha$ is a directional algorithm of $T$, $k\in\{1,\ldots,n\}$, and $Y_{(k)}$ is an RDA of $T_{k}$. In Definition 3.2, we give a precise definition of a new algorithm $\alpha[Y_{(k)}/T_{k}]$ by replacing the behavior on $T_{k}$ of $\alpha$ with $Y_{(k)}$. In Theorem 3.8 (2), we show the following.
Let $X$ be a randomized directional algorithm on $T$ in the broad sense. Then there exists an RDA $Z$ on $T$ with the following properties.
$\max_{\xi\in\Omega(T)}\mathrm{cost}(Z,\xi)\leq\max_{\xi\in\Omega(T)}\mathrm{cost}(X,\xi)$
$\max_{\xi\in\Omega_{0}(T)}\mathrm{cost}(Z,\xi)\leq\max_{\xi\in\Omega_{0}(T)}\mathrm{cost}(X,\xi)$
$\max_{\xi\in\Omega_{1}(T)}\mathrm{cost}(Z,\xi)\leq\max_{\xi\in\Omega_{1}(T)}\mathrm{cost}(X,\xi)$
Theorem 3.8 (2) is shown by means of the following three.
  • [SW86] A known result that there exists an RDA that attains both $d_{0}$ and $d_{1}$.
  • A precise evaluation of cost of $\alpha[Y_{(k)}/T_{k}]$.
  • Induction on the height of the tree.
The item 1 is stated in [SW86, Lemma 3.1] without proof. To keep the paper self-contained, we give a proof the item 1. By Theorem 3.8 (2), we show $R_{\mathrm{dir},i}(T)=d_{i}(T)$. In this way, Problem 1 will be solved in Section 3.
Problem 2:   Problem 2 is more difficult to solve. In Section 4, to solve Problem 2, we investigate the case when $A$ in (i) is a deterministic depth-first algorithm $\alpha$. Instead of $A^{\prime}$ in (iii), we elaborately define a randomized directional algorithm in the broad sense $B_{\alpha,\delta}$, which is a key concept in this study. Figure 1 illustrates interdependence among concepts necessary to define $B_{\alpha,\delta}$.
Figure 1: Interdependence of definitions, simplified version
  • $\alpha$ is a given deterministic depth-first algorithm of $T$.
  • $S$ is a good probability distribution on $T$ setting value of root 00.
  • $\delta$ is the behavior of $S$ just under the root. Thus, $\delta$ is a probability distribution on the truth assignments to nodes just under the root.
  • For each $j=1,\dots,n$ and $i=0,1$, $S^{i}_{j}$ is a good probability distribution that sets value of the root of $T_{j}$ to $i$.
  • $S^{\delta}$ is a probability distribution on $T$ made by combining $\delta$ and $S^{i}_{j}$. At the depth-1 level of $T$, $S^{\delta}$ behaves $S$-like. On each $T_{j}$, $S^{\delta}$ behaves $S^{i}_{j}$-like.
  • $Y_{j}$ is a randomized directional algorithm in the broad sense of $T_{j}$ such that $Y_{j}$ behaves well both for $S^{0}_{j}$ and $S^{1}_{j}$.
  • $B_{\alpha,\delta}$ is a randomized directional algorithm in the broad sense of $T$.
    • At the depth-1 level of $T$, $B_{\alpha,\delta}$ behaves similarly to the movement of $\alpha$ with respect to $S^{\delta}$.
    • On each $T_{j}$, $B_{\alpha,\delta}$ behaves $Y_{j}$-like.
The nickname “chimera” for algorithm $B_{\alpha,\delta}$ arises from the fact that $B_{\alpha,\delta}$ is composed of parts from several other algorithms and probability distributions as well as $\alpha$. We show Inequality (4.57):
$\mathrm{cost}(B_{\alpha,\delta},S^{\delta})\leq\mathrm{cost}(\alpha,S^{\delta})$
This inequality is a key to prove $R_{\mathrm{DF},i}(T)=R_{\mathrm{dir},i}(T)$. In this way, Problem 2 will be solved in Section 4. Figure 1 is a simplified version of Figure 2 in subsection 4.1.
In Section 2, we give the basic definitions and list some fundamental facts regarding equilibrium values. In Section 5, we summarize our results and describe some immediate consequences along with those reported in previous studies. In Section 6, we provide rigorous proofs of some facts used in other sections. Looking back at our proof, we can say that Observation 1.8 oversimplifies the facts. It would be an interesting direction of future research to extend our proof to computational models other than AND-OR trees, in particular trees composed of threshold functions or majority functions [CP22].
2 Definitions and Preliminaries
In this section, we will define some basic concepts and see some fundamental facts on AND–OR trees. See also [AB09, Chapter 12] and [S17] for helpful examples.
Saks and Wigderson [SW86] defined RDA in order to get a recurrence formula of randomized complexity. In general, a probability distribution on the deterministic directional algorithms is not necessarily an RDA .
Example 2.1.